Fill in the missing fractions.
 ▯ + 5 / 27 = 12 / 27 ?????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given▯+ 5 / 27 = 12 / 27
▯= 12 / 27–5 /27
▯= (12–5) / 27
▯= 7 /27

Thanks